# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v2

@@ 3.1 Core-Agent: Keystone component of an LLM-based agent
This issue arises from a lack of modularity in the software design, which can be mitigated by adhering to the Single Responsibility Principle (SRP). As emphasized in the paper [17], the SRP is crucial in software development because it provides granularity at different levels of the software, both in terms of code and functionalities. This granularity facilitates the implementation of future improvements and enhances reusability opportunities. Additionally, applying this principle prevents various code smells, ensuring the modularity of the code and thus making it easier for practitioners to manage and maintain the software.
Paper [18] attempts to establish boundaries between LLM-based agents, the tools they utilize, and their surrounding environment. Although the proposed LLM-based agent system offers an abstract separation of the agent’s components, it does not clearly delineate them from a software engineering perspective. Understanding the theoretical contributions of each component within a unified agent is valuable; however, from a practical development standpoint, it is essential to identify the location, functionality, and role of each internal component within the agent.
Upon a thorough analysis of LLM-powered agents from a software perspective, we recognize an LLM-based agent as a software system comprising various components including tools, LLMs, and the core-agent, our newly introduced term. While the core-agent is not a newly invented component, it serves as a label to denote an existing functional element that has been previously implicit or unnamed in past frameworks or architectures. Essentially, this label highlights its crucial role within these systems. This terminology helps clarify the structure of an LLM-based agent by identifying its essential parts. As depicted in Figure 2, the core-agent interacts with the environment, collaborates with the language model to make decisions, and translates high-level goals into concrete actions optionally by leveraging the available tools.
Core-Agent:
In this architecture, the core-agent serves as the keystone component, acting as the crucial interface between itself and all the other elements of the agent. It facilitates communication and coordination among these various parts to ensure optimal functionality of the entire system. Functioning as the executor of the agent, the core-agent translates plans developed by the Large Language Model (LLM) into actionable steps potentially leveraging tools and engages with the environment to provide the agent with insights into the external world. These capabilities enable the core-agent to operate as a controller, ensuring precise and complex synchronization of actions that result in flawless interactions and effective information sharing.
The significance of the core-agent is underscored by its symbiotic relationship with the LLM, with each enhancing the other’s capabilities. LLMs unveils powerful language understanding, cognitive and reasoning abilities, alongside extensive knowledge. However, it lacks the perception and action components, which are provided by the core-agent, enabling direct environmental interaction. This collaboration expands the scope of problems the agent can address, effectively combining the strengths of both the LLM and the core-agent to tackle a broader range of challenges.
Furthermore, the presence of the core-agent as a controller mitigates the overall complexity of the agent, especially in a multi-LLM setup where it serves as the primary communication hub within the agent. Consequently, information flow between the LLMs is managed through the core-agent, streamlining interactions and ensuring efficient coordination. In terms of communication, the core-agent is capable of interacting with humans through a well-structured pipeline. However, there are instances where the core-agent may not engage in direct communication with humans. This variability is dependent on the specific type of core-agent, as elucidated in detail in Section 3.3.
LLM:
The LLM acts as a cerebral entity covering cognitive tasks such as natural language understanding and comprehensive text generation based on specific context. Indeed, LLMs can communicate with core-agents which in turn can interact with tools or data sources within the system. Considering that an agent can manifest as unimodal or multimodal [18], the latter can be replicated with different task-specific LLMs, as illustrated in Figure 2. This distinction fosters modularity within our LLM-based agent unified modeling framework, making them open for extension and addition of further domain-specific models without having to introduce changes to the overall system architecture, adhering more to the Open-Close principle (OCP) [19, 20].
Tools:
In LLM-based agents, tools are important assets. They can take various forms such as supplementary systems, software applications or even physical devices [12] that extend and enhance the capabilities of an agent. They span across a spectrum of complexity, ranging from basic API integration to sophisticated auxiliary systems designed for specific tasks. In this context, tools can be considered external if they are independent systems which achieve a complete objective, or internal if they cooperate with the core-agent to achieve tasks in the scope of the agent goal.
In conclusion, our proposed terminology for an LLM-based agent, which includes a core-agent as its central coordinating component, serves to enhance clarity and consistency in describing these systems from a software perspective. The underscored significance of the core-agent highlights the importance of investigating its internal structure. Section 3.2 focuses on defining the various modules within the core-agent and how they collaborate to facilitate cognitive tasks, decision-making, and action execution leveraging available tools.
@@ 3.2 Modeling the Core-Agent Internal Structure
The human brain exhibits remarkable modularity, with distinct regions and functionalities working in coordination to facilitate cognitive processes, decision-making, and behavior. Inspired by this organized design, our proposed framework aims to emulate the brain’s modular structure thought the incorporation of five internal modules: the planning module, the memory module, the profile module, the action module and the security module (Figure 3). By integrating the latter concepts, the core-agent serves as the central coordinator, controlling the interactions and information flow between different components, as elaborated in the previous section.
This modular framework effectively addresses the challenges of extendibility and maintainability. Specifically, it allows for independent development and integration of new modules without impacting the entire system. In fact, such modules can be replaced or upgraded separately, facilitating the addition of new capabilities and enabling the framework to adapt more easily to new requirements or technologies. Furthermore, the modular structure promotes code reusability, as individual modules can be shared across different agents or applications, thereby reducing duplication and enhancing consistency.
To provide a robust comparison and underscore the novelty of our framework, it is instructive to examine it alongside existing approaches. A recent survey [5] introduced a framework for LLM-based agents, comprising four key modules: profile, memory, planning, and action modules. For a comprehensive overview of this framework, we direct readers to Section 2. While the survey’s proposed framework treated the LLM-based agent as a whole system without distinguishing a core component from other entities, our adaptation, tailored specifically for the core-agent, necessitates modifying certain definitions and module structures to accommodate the separation between the LLM and the core-agent. Notably, we had to reposition the four key modules to operate under the core-agent within the framework alongside our newly introduced security-oriented module.
@@ 3.2.1 Planning Module
In our proposed framework, the planning module is a pivotal element that enables the agent to break down the complex problems to generate effective plans, the same as in the original framework [5]. However, in our solution the planning module becomes part of the core-agent and collaborates with the other sibling modules to empower the agent to achieve specific goals.
The planning module requires complicated understanding and reasoning [21] which could leverage the capabilities of an LLM. In fact, the LLM’s ability to comprehend nuanced instructions, interpret implicit information, and adapt to various problem domains renders it an invaluable asset in the planning process. Consequently, the planning module can formulate more comprehensive, context-aware, and adaptable plans, significantly enhancing the core-agent’s decision-making process. Furthermore, the planning module works in close collaboration with all other modules within the core-agent, including the memory module for Memory-augmented Planning [22]. This collaboration enhances planning capabilities by leveraging stored information such as commonsense knowledge, past experiences, and domain-specific knowledge.
This section will detail the functionalities and characteristics of this module from four critical aspects: process, strategies, techniques, and feedback sources (Figure 4). These aspects derive from the clear separation we have established between the core-agent and the LLM, and between our system and external systems (Figure 3).
@@
While generating the procedures to undertake, the planning module follows an incremental approach. Therein, the steps dictate how tasks are decomposed, how the planning procedure unfolds, and how alternative solutions are generated and evaluated. Inspired by the human capacity to decompose complex tasks into simpler ones to achieve overarching goals, this process comprises two main steps [22].
@@
To guide the planification, organization and execution of complex tasks, the planning module must adhere to a specific strategy when elaborating the procedures during the plan generation step. Selecting a planning strategy can substantially influence the effectiveness, efficiency, and robustness of the resulting plans. Within the LLM-Agent-UMF, we define two primary strategies:
@@
    Consequently, in single path strategy, any error in one procedure can lead the subsequent procedures or the overall plan to be suboptimal or infeasible [22], negatively impacting the entire strategy. A straightforward approach to mitigate such failures is the Multi-path strategy, which involves two major steps: The first phase involves leveraging the LLM to generate multiple plans for the complex task. Indeed, each intermediate step holds multiple subsequent paths [5]. As for the second phase, it deals with the evaluation and the selection of the most suitable path.
@@
The planning module follows planning techniques as methodological approaches to form executable plans. These techniques are chosen based on criteria such as the complexity of the task, and the need for contextual comprehension. Our framework presents two primary techniques:
@@
    Language Model powered (LM-powered) methodologies leverage the vast knowledge and reasoning capabilities inherent in LMs to orchestrate planning strategies. Within our framework, this category also encompasses neural planners [22], who are adept at addressing intricate and vague tasks necessitating nuanced comprehension and adaptive problem-solving abilities.
This categorization in our framework provides a clear distinction between approaches primarily driven by LMs and those relying on rule-based methods. It allows for a more streamlined understanding of planning techniques within the context of our core-agent architecture, while still acknowledging the valuable contributions of various planning approaches in [5] and [22].
@@
The planning module’s effectiveness is further enhanced by incorporating feedback from various sources, including humans, tools, and sibling core-agents. This feedback loop allows for continuous refinement and adaptation of plans, ensuring that the agent remains responsive and efficient. Moreover, incorporating planning strategies and feedback mechanisms enhances adaptability, enabling the core-agent to address a wide range of scenarios, from simple tasks to complex challenges. Such strategic diversity equips the system with greater intelligence and versatility.

### v3

@@ 3.4 Contributions: Core-Agent as Keystone Component
The core-agent is not a newly invented component but a label to denote an existing functional element that has been previously implicit or unnamed in past frameworks or architectures. It acts as the pivotal interface, orchestrating interactions among Large Language Models (LLMs), tools, and the environment to translate high-level goals into actionable outcomes, enhancing clarity and reusability [30]. By providing a distinct, modular entity, the core-agent mitigates the complexity of agent architectures, particularly in multi-LLM setups, serving as the primary communication hub. Its internal structure, comprising five modules—planning, memory, profile, action, and security—operationalizes this role, as detailed in Section 4 and illustrated in Figure 3. This section outlines the framework’s key contributions, mapping each to the design principles from Section 3.3 that enable actionable concepts to ensure the quality requirements from Section 3.2, such as Modularity, Security, and Scalability.
The framework’s contributions are as follows:
  • Novel Core-Agent Terminology: Labeling the central coordinating component as the “core-agent” addresses terminological discrepancies in LLM-based agent research, fostering consistent communication. This contribution is driven by Terminological Consistency to resolve ambiguities, Modularity, SRP, and Separation of Concerns to delineate the core-agent as a distinct entity, and Constraint-based Design to establish a standardized term for consistent framework behavior, supporting Maintainability and Modifiability [30].
  • Dedicated Security Module: Introducing a security module ensures ethical and secure operations through layered safeguarding mechanisms, addressing critical gaps in prior frameworks that neglected security. This contribution is guided by Security by Design, which embeds robust protection from the outset, Privacy by Design, which prioritizes sensitive data protection, C.I.A. Triad, which ensures Confidentiality, Integrity, and Availability, Trust Boundaries Definition, which secures external interactions, and Defense in Depth, which provides layered Resilience. Additionally, Modularity, SRP, and Separation of Concerns isolate security functions, and Constraint-based Design enforces security constraints, enhancing Trustworthiness, Resilience, and Privacy [13, 28, 27].
  • Enhanced Module Delineation: Refining the core-agent’s modules with new methods and perspectives (e.g., four-aspect planning analysis, three-perspective memory analysis) improves Flexibility, Interoperability, and Performance. This contribution is enabled by Modularity, SRP, and Separation of Concerns for clear module responsibilities, Open-Closed Principle and Loose Coupling for independent Extensibility, Variability Management for diverse module methods, Extensibility Mechanisms for adding new methods, and Constraint-based Design for consistent module behavior, as detailed in Section 4.
  • Active and Passive Core-Agent Classification: Classifying core-agents as active (authoritative, encompassing all five modules for complex task management) or passive (non-authoritative, limited to action and security modules for simplified execution) provides a novel taxonomy to address architectural ambiguities in prior agent designs. This classification enhances role clarity, enabling active core-agents to handle planning and decision-making while passive core-agents streamline task execution with minimal synchronization needs. It is driven by Modularity, SRP, and Separation of Concerns for distinct role delineation, Composability for combining core-agents in multi-core setups, Variability Management for flexible role assignments, and Constraint-based Design for predictable interactions, supporting Scalability and Modularity [30].
  • Multi-Core Agent Architectures: Proposing uniform (multi-passive or multi-active) and hybrid (one-active-many-passive and many-active-many-passive) multi-core architectures addresses Scalability and Interoperability, challenges of multi-agent systems. The one-active-many-passive architecture is optimal, balancing the simplicity of passive core-agents with the managerial capabilities of an active core-agent to dynamically allocate resources and handle diverse tasks. This contribution is enabled by Composability for seamless core-agent integration, Open-Closed Principle and Loose Coupling for extensible and scalable designs, Technology-Agnostic Design for supporting diverse LLMs and tools, Variability Management for flexible configurations, Extensibility Mechanisms for accommodating system growth, and Constraint-based Design for consistent Performance, ensuring Scalability, Flexibility, and robustness [37].
3.5 Comparison with Prior Frameworks
LLM-Agent-UMF introduces a modular architecture centered on the core-agent, comprising five internal modules: planning, memory, profile, action, and security, designed with a technology-agnostic approach and principles like Modularity, Single Responsibility Principle (SRP), Separation of Concerns, Open-Closed Principle, Loose Coupling, and Constraint-based Design to ensure Flexibility, Extensibility, and Maintainability across all modules. This section compares our framework primarily with the prior framework proposed by [5], which defines LLM-based agents from a functional perspective through four modules: planning, profile, memory, and action. We occasionally reference differences with [7] to highlight additional distinctions, without focusing on it. Our framework addresses limitations in [5], such as functional overlaps, ambiguous terminology, and lack of architectural delineation, as illustrated in [5]’s Figure 1, which does not clarify component boundaries, and introduces novelties, such as the security module, to ensure modularity, clarity, and trustworthiness.
  • Planning Module: In [5], the planning module guides the agent in decomposing tasks into subtasks, leveraging past behaviors for responsiveness. Our framework aligns with this functional perspective but repositions the planning module within the core-agent, enabling collaboration with sibling modules, such as memory for memory-augmented planning [19]. We introduce a novel four-aspect analysis (process, strategies, techniques, feedback sources) to model the aspects of planning from multiple perspectives, as opposed to [5]’s abstract functional description. This structured approach mitigates functional overlaps, particularly with memory, and clarifies planning responsibilities, supporting Modularity and Composability. Similarly, [7]’s “Rethinking” module, functionally akin to planning, guides future actions through introspection on past actions and feedback, but its separation as a distinct module and lack of a core-agent structure limit architectural clarity compared to our integrated approach. This module’s design was further guided by Extensibility Mechanisms to support flexible planning strategies and techniques.
  • Memory Module: [5] defines the memory module as retaining internal logs of thoughts, actions, and observations, using ambiguous terms like Unified Memory (short-term, in-context learning) and Hybrid Memory (short- and long-term). This abstract approach leads to functional overlaps, particularly with reflection on the memory, which [5] includes in the memory module. Our framework redefines memory through a three-perspective analysis (scope, location, format) to manage storage and retrieval of task-related information, building on [7]’s short-term and long-term categorization but enhancing it with alignment to established human memory models [18] and moving all reflection aspects to the planning module. This adaptation eliminates overlaps, resolves ambiguous terminology, and enhances architectural clarity, ensuring Maintainability. Similarly, [7] treats memory as a state repository, with reflection handled by its “Rethinking” module, functionally akin to planning, aligning with our functional approach, but its lack of a core-agent structure and structured analysis reduces clarity compared to our modular design.
  • Profile Module: [5]’s profile module delineates the LLM’s diverse roles, supporting three methods: Handcrafted In-Context Learning, LLM-Generation, and Dataset Alignment. Our framework builds on [5]’s role-defining function, refining the conceptual and functional clarity of its three methods (Handcrafted In-Context Learning, LLM-Generation, Dataset Alignment) to support Dynamic Adaptation, and introduces a novel fourth method, Fine-tuned Pluggable Modules, which enhances Performance by reducing context size and memory footprint for efficient inference. These enhanced and new methods, integrated within the core-agent, ensure clear separation from the LLM, addressing [5]’s lack of architectural boundaries and improving Flexibility and Performance. Unlike [5], [7] does not define a distinct profile module, relying on LLM configuration within its quintuple model, which limits role-specific adaptability compared to our framework. This module was further shaped by Variability Management and Extensibility Mechanisms to enable diverse and pluggable role-defining methods.
  • Action Module: In [5], the action module translates decisions into outputs using external tools, influenced by planning, memory, and profile modules. Our framework retains this functional role, translating plans into executable actions guided by specific triggers and goals, but refines it with a four-aspect analysis (goal, trigger, action space, impact), introducing new triggers (Plan Following and API Call Request) to enhance action initiation clarity. Unlike [5], which risks overlap between action and memory functions (e.g., Action via Memory Recollection), our framework separates these responsibilities, ensuring the action module focuses on execution while memory handles storage, thus improving Modularity and Efficiency. Similarly, [7] defines action as tool usage or messaging, with its “Rethinking” module, functionally akin to planning, guiding future actions through introspection on past actions and feedback, whereas our framework integrates reflection within the planning module’s feedback sources and action module’s impact analysis for clearer architectural delineation within the core-agent. The action module was further guided by Extensibility Mechanisms to support new triggers and flexible tool integration.
  • Security Module: [5] does not include a security module, leaving gaps in addressing ethical use, privacy, and robustness, critical concerns for LLM-based agents [13]. Our framework introduces a novel security module within the core-agent, ensuring ethical and secure operations through layered safeguarding mechanisms (prompt, response, data privacy). Likewise, [7] omits a dedicated security module, focusing on cognitive and functional aspects without addressing architectural security, further underscoring our framework’s novelty. This module’s design was driven by Security by Design, Privacy by Design, C.I.A. Triad, Trust Boundaries Definition, and Defense in Depth to ensure robust protection against ethical and security risks.
By adapting [5]’s modules to operate under the core-agent, introducing a security module, and addressing deficiencies also present in [7], the LLM-Agent-UMF overcomes functional overlaps, ambiguous terminology, and lack of architectural clarity.
3.6 Validation Approach
The validation of the LLM-Agent-UMF was conducted through a scenario-driven approach, AFTRAM (Section 3), strategically designed to assess its capability to guide the architectural design of LLM-based agents without implementing them. The goal was to evaluate the framework’s effectiveness in enhancing Modularity, Scalability, Reusability, Maintainability, Flexibility, Security, and Privacy, as outlined in the quality requirements (Section 3.2). This was achieved by applying the framework to thirteen state-of-the-art agents and designing five novel multi-core agent architectures, each targeting specific quality attributes. The agents were carefully selected to represent a diverse range of functionalities, including tool usage (Toolformer [38], Confucius [39], ToolAlpaca [40], Gorilla [41], ToolLLM [42], GPT4Tools [43]), memory management (ChatDB [23]), planning (Chameleon [8], LLM+P [44]), chemistry applications (ChemCrow [12]), security (LLMSafeGuard [45]), and proprietary systems (ChatGPT 4o and ChatGPT 4o mini). This diversity ensured comprehensive testing of the framework’s applicability across varied domains and complexities.
The validation began with assessing the framework’s ability to clarify architectural components by identifying implicit core-agents in existing systems (Section 5.1). For instance, Toolformer and ToolLLM, which do not explicitly define themselves as agents, were analyzed to uncover core-agents, focusing on Modularity, Reusability, and Terminological Consistency. This process confirmed the framework’s capacity to delineate components clearly, enhancing Maintainability and Interoperability. Next, the internal structure of core-agents was evaluated by classifying them as active or passive across the thirteen agents (Section 5.2). This analysis targeted Modularity, Flexibility, and Security, revealing gaps, such as the absence of security modules in 78% of tool-using agents (e.g., ToolAlpaca, Gorilla), except in ChemCrow and ChatGPT 4o, which demonstrated robust privacy measures. The classification also highlighted Scalability, as passive core-agents (e.g., in Toolformer, Confucius) required less synchronization, unlike active core-agents (e.g., in Chameleon, LLM+P).
The framework’s strength in designing scalable and flexible multi-core architectures was tested through five novel agent designs (Section 5.3). First, LLM-based Agent 1 (LA1) combined Toolformer and Confucius as a multi-passive core-agent system, assessing Scalability, Reusability, and Flexibility. This design leveraged their complementary tool-handling capabilities, but the framework identified a privacy gap in data transfers, underscoring its ability to pinpoint security needs. Second, LLM-based Agent 2 (LA2) explored ToolLLM and ChatDB integration in two variants: LA2-A (multi-active core-agents) and LA2-B (monolithic active core-agent). This scenario evaluated Scalability, Modularity, and Maintainability, addressing synchronization challenges in LA2-A and potential module conflicts in LA2-B, demonstrating the framework’s guidance in resolving architectural complexities. Third, LLM-based Agent 3 (LA3) integrated LLMSafeGuard’s security module into ToolLLM, focusing on Security, Privacy, and Modifiability. The seamless incorporation highlighted the framework’s support for modular extensions without functional conflicts. Finally, LLM-based Agent 4 (LA4) proposed a hybrid one-active-many-passive architecture, combining LLM+P’s active core-agent, LLMSafeGuard’s security module, and Toolformer/Confucius passive core-agents. This design assessed Scalability, Flexibility, Security, and robustness, leveraging LLM+P’s advanced planning and addressing privacy vulnerabilities in tool usage. By strategically selecting agents with diverse capabilities and designing these scenarios, the LLM-Agent-UMF proved its capacity to guide the creation of modular, secure, and scalable agent architectures, identifying risks and opportunities without requiring implementation.
4 The LLM-Agent-UMF
The LLM-based Agent Unified Modeling Framework provides a modular and scalable architecture for designing LLM-based agents, addressing limitations in existing frameworks by introducing the core-agent as a central coordinator. This section details the framework’s components, including the core-agent’s internal modules, active/passive classifications, and multi-core architectures, emphasizing Modularity, Security, and Flexibility.
4.1 Overview
The LLM-Agent-UMF redefines LLM-based agent design by introducing the core-agent as the keystone component, orchestrating interactions among Large Language Models (LLMs), tools, and the environment to achieve high-level goals. Inspired by the human brain’s modularity, the core-agent comprises five modules—planning, memory, profile, action, and security—designed for Extensibility and Maintainability, as shown in Figure 3. The framework classifies core-agents as active (authoritative, managing complex tasks) or passive (non-authoritative, executing specific actions), enabling versatile architectures like the hybrid one-active-many-passive setup, which balances simplicity and functionality. By adhering to software engineering principles such as the Single Responsibility Principle and Open-Closed Principle, LLM-Agent-UMF ensures clarity, reusability, and robust security, overcoming the architectural ambiguities and functional overlaps of prior frameworks [5, 7].
@@ 4.2 Modeling the Core-Agent Internal Structure
Inspired by the human brain’s modularity, our framework incorporates five internal modules within the core-agent: planning, memory, profile, action, and security (Figure 3). The core-agent acts as the central coordinator, managing interactions and information flow among these modules, as described in the previous section. This modular structure enables independent development and replacement of modules, supporting the addition of new capabilities and adaptation to new technologies. It also facilitates code reusability, allowing modules to be shared across different agents or applications.
@@ 4.2.1 Planning Module
The planning module is a pivotal component of the core-agent, enabling the agent to decompose complex problems and generate effective plans. It leverages the LLM’s capabilities for understanding nuanced instructions, interpreting implicit information, and adapting to diverse problem domains, producing comprehensive, context-aware plans that enhance the core-agent’s decision-making. The module collaborates closely with other core-agent modules, such as the memory module for memory-augmented planning [19], utilizing stored information like commonsense knowledge and past experiences. This section details the module’s functionalities across four aspects: process, strategies, techniques, and feedback sources (Figure 4).
@@
The planning module adopts an incremental approach to generate procedures to undertake, systematically decomposing tasks, unfolding the planning process, and generating and evaluating alternative solutions. This process comprises two main steps [19]: task decomposition and plan generation.
@@
To guide the planning and execution of complex tasks, the planning module adheres to a specific strategy during the plan generation step. Selecting a strategy influences the effectiveness, efficiency, and robustness of the resulting plans. Within the LLM-Agent-UMF, we define two primary strategies:
@@
    Consequently, in the single-path strategy, an error in one procedure can render subsequent procedures or the overall plan suboptimal or infeasible [19]. The multi-path strategy mitigates such failures through two steps: First, the LLM generates multiple plans for the complex task, with each intermediate step offering multiple subsequent paths. Second, the module evaluates and selects the most suitable path.
@@
The planning module follows planning techniques as methodological approaches to construct executable plans. These techniques are chosen based on criteria such as the complexity of the task, and the need for contextual comprehension. Our framework presents two primary techniques:
@@
    Language Model powered (LM-powered) methodologies leverage the vast knowledge and reasoning capabilities inherent in LMs to orchestrate planning strategies. Within our framework, this category also encompasses neural planners [19], adept at addressing intricate and vague tasks requiring nuanced comprehension and adaptive problem-solving.
This categorization distinguishes LM-driven approaches from rule-based methods, streamlining the understanding of planning techniques within the core-agent architecture.
@@
The planning module’s effectiveness is further enhanced by incorporating feedback from various sources, including humans, tools, and sibling core-agents. This feedback loop allows for continuous refinement and adaptation of plans, ensuring that the agent remains responsive and efficient. Moreover, incorporating planning strategies and feedback mechanisms enhances Flexibility, enabling the core-agent to address a wide range of scenarios, from simple tasks to complex challenges. Such strategic diversity equips the system with greater intelligence and versatility.